# LLM-Enhanced Static Analysis for Precise Identification of Vulnerable OSS Versions

### v1

@@ 5.5. Threats to Validity
6. Discussion
Time Efficiency. In our experimental setup, we took 10 days to build the software and version pool. Then we took 30 days to build the ground-truth dataset, including the time taken to establish the dependent environment for each software version, as well as the time to perform vulnerability verifying and labeling. Vercation analyzes each vulnerability on an average of 28.61 seconds.
Branch Fixing Oversights. In some cases, we observe that security fixes are promptly incorporated into the latest version branches of OSS. However, rectifying oversights frequently transpires during the backporting of fixes to older release branches. In such scenarios, projects not utilizing the latest branch version may encounter a deficiency in receiving security updates. Simultaneously, the latest versions updated in older branches remain exposed to risks.
Limitation and Future Work. Vercation operates under certain assumptions that may limit its applicability. Firstly, it is tailored for C/C++ projects, yet its methodology can be applied to other programming languages. With an accurate exploration of common vulnerability types in diverse languages, Vercation could identify vulnerable versions in those languages. Secondly, Vercation exclusively addresses vulnerabilities within functions, thus omitting inter-functional vulnerabilities. However, we are contemplating the integration inter-functional data-flow information into the vulnerability signatures to tackle such scenarios. In the evaluation results presented in section 5.2, Vercation demonstrated notable precision (93%) and recall (98%). There exists room for further enhancement in both precision and recall. In semantic similarity comparisons, the utilization of heavyweight program analysis techniques, such as symbolic execution, could enhance the understanding of code semantics, thereby reducing false negatives.
7. Related work
@@
Vulnerability Fixing Commit Analysis. Several studies have focused on analyzing vulnerability fixing commit (VFC) to extract vulnerability features, which can be used for various purposes such as vulnerability detection, classification, and affected version identification. VulPecker (Li et al., 2016) extracts code patterns from VFC to detect similar vulnerabilities in other software versions. SySeVR (Li et al., 2021) uses a systematic approach to extract syntax- and semantic-based vulnerability features from patches. VulDeePecker (Li et al., 2018) leverages deep learning techniques to learn vulnerability patterns from code gadgets extracted from vulnerability fixing commits. VulCNN (Wu et al., 2022) uses convolutional neural networks to learn vulnerability patterns from code changes in VFC. Most existing VFC analysis methods rely heavily on predefined rules or patterns, which may not capture the full semantic context of vulnerabilities. They also often struggle with handling complex code changes and may miss subtle vulnerability indicators. Vercation incorporates advanced VFC analysis by leveraging LLMs to understand the semantic context of code changes in fixing commits, potentially capturing more nuanced vulnerability features.
@@
Vulnerable Detection for Open-source Software. Several approaches have been proposed for vulnerable version identification. V-SZZ (Bao et al., 2022) uses the SZZ algorithm to backtrace patch commits and identify vulnerability-introducing changes. MVP (Xiao et al., 2020) employs program slicing to extract vulnerability and patch signatures, then uses these to identify potentially vulnerable functions. AFV (Shi et al., 2022) introduces a vulnerability-centric approach for precise analysis of affected versions in web vulnerabilities.V0Finder (Woo et al., 2021) generates fingerprints for vulnerable functions and uses a clone-based technique to detect vulnerable versions across different software releases. VCCFinder (Perl et al., 2015) uses machine learning techniques to identify vulnerability-contributing commits. VUDDY (Kim et al., 2017) proposes a scalable approach for vulnerable code clone detection in large-scale code bases. Existing approaches often rely on predefined patterns and syntactic-level analysis, leading to imprecision in vulnerability characterization and code clone detection. Vercation combines static analysis with LLM for semantic understanding and introduces AST-based semantic-level code clone detection.
@@
Code Refactoring Detection. Some researchers conducted studies on the characteristics of code refactoring not altering the semantic behavior of the code and have proposed detection methods. MLRefScanner (Noei et al., 2024) detects refactoring commits in machine-learning Python projects by analyzing commit history and extracting features that represent refactoring activities without altering semantic behavior. Abid et al. (Abid et al., 2020) perform a study outlining different levels of refactoring from code-level to architecture and discuss the importance of maintaining behavior during refactoring. Eman et al. (AlOmar et al., 2021) explores refactoring practices in industrial settings, focusing on how refactoring is integrated into the code review process while ensuring that the original software behavior is preserved. Existing code refactoring detection methods primarily focus on high-level changes like renaming, method/class moves. Our approach, however, targets sentence-level changes through fine-grained AST expansion and normalization, including inline function expansion during AST generation. This level of detail in internal code structure analysis is a novel aspect not previously considered.
8. Conclusion
@@
This paper proposed Vercation, an approach designed for vulnerable version identification of open-source C/C++ software. The core idea involves a symbiotic combination of static analysis, LLM, and semantic-level code clone detection. Our approach introduces two key innovations: leveraging LLMs to enhance extraction of vulnerability-related statements and employing semantic-level code clone detection based on expanded Abstract Syntax Trees (ASTs). Experimental results on our dataset of 74 CVEs across 11 popular open-source projects validate that Vercation surpasses existing techniques for vulnerable version identification, demonstrating notable improvements in both precision and recall. Equipped with the discovery from Vercation, developers can effectively address potential threats arising from cross-version vulnerabilities in OSS, thereby enhancing the overall security of the software ecosystem.

### v2

@@ VI-C1 Internal Validity
VI-C2 External Validity
Language generalisability. Vercation’s core methodology is language-agnostic and can be adapted to other programming languages. The program slicing tool Joern inherently supports multiple languages, including Java, Python, JavaScript, and PHP. The main adaptations required would be updating the sensitive function table for language-specific vulnerability patterns and modifying AST generation rules according to the target language’s syntax and semantics. Future work could involve extending Vercation to support a broader range of programming languages and verifying its effectiveness across different language paradigms.
Repository branching complexity. While our approach handles common merge [70] and branching patterns effectively, complex branching scenarios, such as when the same vulnerable code is independently modified in multiple parallel feature branches before being merged (creating diamond-shaped merge patterns), may require more sophisticated graph traversal algorithms to accurately determine the earliest vulnerability introduction point [71, 72].
VII Related work
VII-A Vulnerability Fixing Commit Analysis
@@
Several studies have focused on analyzing vulnerability fixing commit (VFC) to extract vulnerability features, which can be used for various purposes such as vulnerability detection, classification, and affected version identification. VulPecker [73] extracts code patterns from VFC to detect similar vulnerabilities in other software versions. SySeVR [74] uses a systematic approach to extract syntax- and semantic-based vulnerability features from patches. VulDeePecker [75] leverages deep learning techniques to learn vulnerability patterns from code gadgets extracted from vulnerability fixing commits. Most existing VFC analysis methods rely heavily on predefined rules or patterns, which may not capture the full context of vulnerabilities. Vercation incorporates VFC analysis by leveraging LLMs to understand the context of code changes in fixing commits, potentially capturing more nuanced vulnerability features.
VII-B Vulnerable Version Detection for OSS
@@
Several approaches have been proposed for vulnerable version identification. V-SZZ [17] uses the SZZ algorithm to backtrack vfcs and identify vulnerability-introducing changes. MVP [12] employs program slicing to extract vulnerability and patch signatures, then uses these to identify potentially vulnerable functions. V0Finder [16] generates fingerprints for vulnerable functions and uses a clone-based technique to detect vulnerable versions across different software releases. VCCFinder [76] uses machine learning techniques to identify vulnerability-contributing commits. VUDDY [77] proposes a scalable approach for vulnerable code clone detection in large-scale code bases. Existing approaches often rely on predefined patterns and syntactic-level analysis, leading to imprecision in vulnerability characterization and code clone detection. Vercation combines static analysis with LLM for code understanding and introduces AST-based code clone detection to address the code structure modification problem.
VII-C Code Refactoring Detection
@@
Some researchers conducted studies on the characteristics of code refactoring not altering the code behavior of the code and have proposed detection methods. MLRefScanner [78] detects refactoring commits in Python projects by analyzing commit history and extracting features that represent refactoring activities without altering code behavior. Abid et al. [79] perform a study outlining different levels of refactoring from code level to architecture and discuss the importance of maintaining behavior during refactoring. Eman et al. [80] focus on how refactoring is integrated into the code review process while ensuring that the original software behavior is preserved. Existing code refactoring detection methods primarily focus on high-level changes like renaming and method/class moves. Our approach, however, targets statement-level changes through fine-grained AST expansion and normalization, including inline function expansion during AST generation.
VII-D Code Clone Detection
Some SOTA techniques utilize static analysis to conduct code clone detection. SCDetector [81] feeds tokens with graph detail into a Siamese architecture neural network to train a code clone detector. Fang et al. [82] propose a joint code representation that applies fusion embedding techniques to learn hidden features of source codes, then train a supervised deep learning model to detect functional code clones. FCDetection [67] generates AST and CFG representations and uses Word2vec to embed these features into vectors. Then the fused feature vectors are input into a deep neural network model for classification of code clones.
VIII Conclusion and Future Work
@@
This paper proposed Vercation, an approach designed for vulnerable version identification of open-source C/C++ software. Our approach introduces two key innovations: leveraging LLMs to enhance the extraction of vulnerability-related statements and employing code clone detection based on expanded and normalized ASTs. Experimental results on our dataset of 122 CVEs across 12 popular open-source projects validate that Vercation surpasses existing techniques for vulnerable version identification, demonstrating notable improvements in both precision and recall. Vercation’s novelty lies not in individual techniques but in the systematic integration of semantic analysis and methodological innovations in LLM application. The combination creates a qualitatively different approach that advances the state-of-the-art in both accuracy and practical applicability, opening new research directions for AI-assisted software security analysis.
In the future, we plan to empirically validate Vercation on a broader set of languages (Java, Python, JavaScript, and PHP) to confirm its effectiveness. Furthermore, we are interested in exploring integrating inter-procedural analysis techniques to enhance the detection of cross-functional vulnerability patterns.
Acknowledgement
We thank the anonymous reviewers for their insightful comments. This research is supported by Beijing Natural Science Foundation (Grant No. L234033), the National Research Foundation, Singapore, and the Smart Nation Group under the Smart Nation Group’s Translational R&D Grant (Award No. TRANS2023-TGC02). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore or the Smart Nation Group.